Q4. You have 600 small marbles in a bag which are Yellow or Red colors. You know that 20% of the marbles are the color Red.
Answer #1

Let X be the number of red marbles picked from the bag.

X\sim Bin(n,p)

P.M.F of X

P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}

total number of marbles to be drawn: n=5

Probability of picking a red marble : p=20/100=0.2

P(X=1)=\binom{5}{1}0.2^1(1-0.2)^{5-1}

=\binom{5}{1}*0.2^1*0.8^4=5*0.2*0.4096=0.4096

Probability that the family member decides where to order take out dinner tonight is 0.4096